How to do calculation 223.13 = x + 5%. What is x? I don't know how to do this.
Answer by stanbon(75887)

223.13 = x + 5%. What is x
223.13 = x + 0.05
x = 223.13 - 0.05
x = 223.08
===============
Cheers,
Stan H.